An appliance rated at 55W is plugged into a 115 V line. How much current is drawn? and if the appliance was an ohmic device, what would be its resistance in (omega)
Answer #1

We know that power P=VI

Here P=55w, V=115v

I = P/V =55/115

I =0.478A. R=V/I =115/0.478 =240.58 ohm